OpenStudy (anonymous):

Let f(x)=3x+1 and g(x)=2-x. Determine values for a such that f(a^2)=g(2a). How do you solve this? Please help D:

OpenStudy (anonymous):

in f(x) substute a^2 for x in g(x) use 2a in place of x you are checking to see if you get the same values on both sides

OpenStudy (radar):

\[3a ^{2}+1=2-2a\] \[3a ^{2}+2a-1=0\] (a+1)(3a-1)=0 a=-1 a=1/3

OpenStudy (anonymous):

Ok well, since you're trying to find where f(a^2) = f(2a), we have to figure out what f(a^2) is and what g(2a) is. f(a^2) = 3(a^2) + 1 f(a^2) = 3a^2 + 1 now let's plug in 2a for g(x): g(2a) = 2 - 2a Ok since, you need to find when f(a^2) = g(2a) just set them equal to each other: 3a^2 + 1 = 2 -2a 3a^2 + 2a + 1 - 2 = 0 3a^2 + 2a - 1 = 0 If you factor that, you get: (3a + 1) (a - 1) = 0 therefore a=(1/3), (1)

OpenStudy (anonymous):

sorry i meant (3a - 1) (a+1) which means it should be: a = (1/3), (-1)

OpenStudy (radar):

Did you follow b100?

OpenStudy (anonymous):

Yes that's the answer in the textbook! Thanks everyone! :)

OpenStudy (radar):

My post follows JuanitaM, do what she suggested, and then my post would be a continuation of the first post. Hermeezey post shows the complete solution starting with what was given in the problem. Just wondering if you follow?

OpenStudy (radar):

O.K. great, good luck in your studies.